# Federated Class-Incremental Learning with Hierarchical Generative Prototypes

### v3

@@ Abstract
Federated Learning (FL) aims at unburdening the training of deep models by distributing computation across multiple devices (clients) while safeguarding data privacy. On top of that, Federated Continual Learning (FCL) also accounts for data distribution evolving over time, mirroring the dynamic nature of real-world environments. While previous studies have identified Catastrophic Forgetting and Client Drift as primary causes of performance degradation in FCL, we shed light on the importance of Incremental Bias and Federated Bias, which cause models to prioritize classes that are recently introduced or locally predominant, respectively. Our proposal constrains both biases in the last layer by efficiently fine-tuning a pre-trained backbone using learnable prompts, resulting in clients that produce less biased representations and more biased classifiers. Therefore, instead of solely relying on parameter aggregation, we leverage generative prototypes to effectively balance the predictions of the global model. Our method significantly improves the current State Of The Art, providing an average increase of $\texttt{+}7.8\%$ in accuracy. Code to reproduce the results is provided in the suppl. material.
@@ 1 Introduction
The traditional paradigm in Deep Learning necessitates accessing large-scale datasets all at once, which hinders scalability and raises significant privacy concerns, especially when sensitive data is involved. Although distributing training across many devices could be an effective solution, there is still no effective mechanism for blending the resulting trained models into a single unified one. Federated Learning (FL) (McMahan et al. 2017) addresses this challenge through a centralized server that coordinates distributed devices, aiming to create a single unified model while minimizing communication costs.
@@
Federated Class-Incremental Learning (FCIL) (Yoon et al. 2021; Dong et al. 2022; Zhang et al. 2023b) takes a step further and couples distributed training with Online Learning, tolerating distribution shifts in the data over time. This presents new challenges, as deep models learning online (without relying on old examples) experience severe performance degradation due to Catastrophic Forgetting (McCloskey and Cohen 1989). In FCIL, the training process unfolds in tasks, each of which shifts the data distribution by introducing new categories. Each task is divided into communication rounds, wherein the local models train on their private data distribution. After local training, each client may transmit information to the orchestrator (server), which creates a global model and redistributes it to all clients. In the literature, some methodologies account for architectural heterogeneity (i.e., heterogeneous FL (Diao, Ding, and Tarokh 2021; Kim et al. 2022b; Ilhan, Su, and Liu 2023)), while others aim to enhance the performance of local models without necessarily converging to a global one (i.e., personalized FL (Collins et al. 2021; Ma et al. 2022; OH, Kim, and Yun 2022)). Instead, we follow the original FCIL setting as presented in (Dong et al. 2022), with the goal of training a single global model in a distributed way.
@@
When incrementally learning on a sequence of tasks, the model struggles the most at differentiating classes from distinct tasks, whereas it works well at separating those within the same one. Albeit one would intuitively link such behavior to Catastrophic Forgetting, it primarily occurs because tasks are learned in isolation, and some classes are never seen simultaneously (Kim et al. 2022a). This causes a bias towards recently introduced classes (Wu et al. 2019), which we refer to as Incremental Bias (IB). The authors of (Luo et al. 2021) observe a similar tendency in the Federated Learning scenario: since clients train exclusively on their local datasets, they exhibit a bias towards their local label distribution. We refer to this effect as Federated Bias (FB).
@@
Figure 1: Federated bias. Histogram of the clients’ responses, computed on the global test set, for $\beta=0.05$. Class indexes have been rearranged for improved visualization (left). The entropy of the response histograms, averaged on all clients, compared with FL performance (right).
@@
In contrast to the well-known Client Drift (Gao et al. 2022; Karimireddy et al. 2020; Zhao et al. 2018), which causes misalignment between the clients’ learned parameters, Federated Bias affects the clients’ responses. Specifically, FB induces clients’ outputs to diverge in different directions, mirroring the patterns of their local label distributions. Also, the strength of FB increases with growing heterogeneity, suggesting a correlation with declining performance (see Section 2). To relieve such an effect, we constrain FB to the last layer by leveraging a frozen pre-trained backbone and efficiently fine-tuning it via prompt learning (Li and Liang 2021). Ideally, prompting keeps the clients’ representations close to the pre-training optimum (hence, close to each other), thus minimizing their Federated Bias. In Section 4.3, we experimentally verify that prompting leads to reduced bias in the feature space w.r.t. fine-tuning all parameters. Consequently, this confines the impact of FB to the last layer, providing the centralized server with less biased representations. On top of that, prompt-based methodologies i) have demonstrated SOTA results in Class-Incremental Learning (Wang et al. 2022b, a; Smith et al. 2023) and ii) adapt only a small portion of the clients’ parameters, enabling unmatched communication efficiency in distributed scenarios (Zhao et al. 2023; Liu et al. 2023).
@@
The authors of (Wu et al. 2019; Zhang et al. 2023a; Luo et al. 2021) address either IB or FB by fine-tuning the classification layer (where such biases are the most evident) on IID data samples. To meet the privacy requirements of FL, which prohibit transferring real data, we follow recent studies (Zhang et al. 2023a; Luo et al. 2021) and leverage latent generative replay. Specifically, at each communication round, we alleviate both biases – previously enforced to the last layer by the adopted prompt-based fine-tuning – by rebalancing the global classifier on a dataset of generated representations. In contrast to other approaches relying on prototypes (i.e., the average feature vectors) to regularize clients’ training procedures (Tan et al. 2022; Guo et al. 2024), we propose to compute their covariance matrix and parameterize a Multivariate Gaussian distribution for each class-client combination. This forms a grid of $num\_classes\times num\_clients$ generative prototypes, which are sampled hierarchically (first by class, then by client) to generate new data points.
@@ 2 On Federated Bias
This section investigates how Federated Bias in clients’ responses is associated with performance degradation in Federated Learning. All experiments are conducted on the CIFAR-100 (Krizhevsky, Hinton et al. 2009) dataset, where the data is heterogeneously distributed across $10$ clients under the commonly adopted distribution-based label imbalance setting (Li et al. 2022; Yurochkin et al. 2019).
To effectively show the presence of FB in the local models, we evaluate them at their most biased state: specifically, at the end of the local training, prior to any synchronization with the centralized server. In Figure 1 (left), we show the histogram of the responses given by two randomly selected clients, trained with $\beta=0.05$, on the global test set. Additionally, we compare them against a model trained conventionally on the global data distribution (referred to as Joint). It can be observed that clients’ predictions are significantly skewed, mirroring their local label distribution.
@@ 3.1 Problem definition
Federated Class-Incremental Learning (Zhang et al. 2023b; Guo et al. 2024) tackles a classification problem across $C$ classes, which are introduced sequentially over $T$ incremental tasks. For each task, the data is distributed in a non-IID manner among $M$ clients. Let $D^{t}$ be the global partition for task $t$, which is split among the $M$ clients, with $D_{m}^{t}$ being the local partition of client $m$ at task $t$. The training procedure of each task is divided into communication rounds, each consisting of a certain number of epochs. At the end of the local optimization, the clients synchronize with the server by exchanging their learnable parameters. The server aggregates these parameters into a global model and redistributes it to all clients, thus concluding the communication round. Since data from previous tasks is unavailable, client $m$ can only train on its dataset $D_{m}^{t}$ during task $t$. Let $f_{\theta_{m}}$ be the local model of client $m$, parameterized by $\theta_{m}$. The local objective is to minimize the loss function $\mathcal{L}$ with respect to the local dataset $D_{m}^{t}$, namely:
@@ 3.2 Hierarchical Generative Prototypes
Differently from other prompt-based CL approaches, which rely on a pool of task-generic (Wang et al. 2022b) or task-specific (Smith et al. 2023; Wang et al. 2022a) prompts, we propose learning a single prompt shared across all tasks. This eliminates the need for prompt selection – which requires an additional forward pass – and significantly accelerates both training and inference. Specifically, we instantiate two learnable vectors, $P_{k}$ and $P_{v}$, for each of the first $5$ transformer blocks, obtaining the prompt $\mathcal{P}=\{P_{k}^{j},P_{v}^{j}\;|\;j\in 1,\ldots,5\}$. Following prefix-tuning (Li and Liang 2021), $P_{k}^{j}$ and $P_{v}^{j}$ are respectively prepended to the keys and values of the $j^{th}$ Multi-head Self Attention layer and optimized jointly to minimize the loss function.
@@
For a given class $c$, the server receives $M$ generative prototypes, one from each client, and seeks to identify the single distribution that best aligns with all of them. To this aim, we take the distribution that minimizes the Jensen-Shannon Divergence (JSD) (Lin 1991). Considering multiple distributions $\mathcal{Q}=\{Q_{1},\ldots,Q_{M}\}$, with importance weights $\Pi=\{\pi_{1},\ldots,\pi_{M}\}$, their JSD is defined as:
@@
Notably, the optimal solution to this problem is to set $\tilde{Q}_{c}$ equal to the distribution $G$. Since, in our case, all distributions $\{Q_{1},\ldots,Q_{M}\}$ are Gaussians, the closest distribution $G$ is precisely defined as a Gaussian Mixture Model (GMM). Following a common practice in Federated Learning (Tan et al. 2022; McMahan et al. 2017), we assign the importance weights $\{\pi_{1,c},\ldots,\pi_{M,c}\}$ to the generative prototypes based on the number of samples from class $c$ observed by client $m$.
@@
Finally, the server redistributes the updated classifier, alongside the aggregated prompt, back to the clients, enabling them to start the next communication round.
Partition $\beta=0.5$ $\beta=0.1$ $\beta=0.05$ $\alpha=6$ $\alpha=4$ $\alpha=2$
Joint $\boldsymbol{92.3}$
EWCFL   $8.7$ $(24.5)$   $4.9$ $(13.1)$   $4.0$ $(11.3)$   $5.4$ $(17.8)$   $4.7$ $(14.4)$ $1.2$ $(6.8)$
LwFFL $16.3$ $(35.8)$   $7.7$ $(17.8)$   $7.7$ $(14.9)$ $11.4$ $(24.6)$ $10.3$ $(19.4)$ $3.0$ $(9.3)$
iCaRLFL $54.8$ $(63.7)$ $48.2$ $(53.5)$ $42.9$ $(43.8)$ $42.2$ $(48.6)$ $38.5$ $(42.3)$ $26.8$ $(28.3)$
CCVR $64.2$ $(79.6)$ $63.2$ $(77.7)$ $61.3$ $(76.4)$ $65.3$ $(79.1)$ $66.4$ $(78.9)$ $51.1$ $(69.7)$
L2PFL $77.4$ $(83.3)$ $71.3$ $(79.4)$ $69.4$ $(77.0)$ $72.8$ $(80.3)$ $70.4$ $(79.2)$ $63.4$ $(72.0)$
CODA-PFL $61.3$ $(71.8)$ $18.1$ $(35.4)$ $15.9$ $(34.3)$ $35.2$ $(50.0)$ $15.1$ $(34.6)$ $27.7$ $(9.5)$
TARGET $13.3$ $(33.8)$   $1.2$ $(13.6)$   $4.0$ $(12.7)$   $6.4$ $(23.9)$   $5.7$ $(18.7)$ $1.6$ $(8.2)$
GLFC -- ()(-)( - ) -- ()(-)( - ) -- ()(-)( - ) $50.0$ $(66.9)$ $50.5$ $(61.0)$   $4.4$ $(12.5)$
LGA -- ()(-)( - ) -- ()(-)( - ) -- ()(-)( - ) $62.9$ $(73.5)$ $61.4$ $(67.5)$   $9.8$ $(15.1)$
PLoRA $79.3$ $(85.9)$ $78.2$ $(84.6)$ $77.8$ $(84.5)$ $79.2$ $(85.3)$ $77.4$ $(83.7)$ $73.2$ $(81.1)$
HGP (ours) $\boldsymbol{89.6}$ $(\boldsymbol{93.1})$ $\boldsymbol{89.7}$ $(\boldsymbol{93.0})$ $\boldsymbol{89.0}$ $(\boldsymbol{92.5})$ $\boldsymbol{89.6}$ $(\boldsymbol{93.3})$ $\boldsymbol{89.7}$ $(\boldsymbol{93.1})$ $\boldsymbol{88.4}$ $(\boldsymbol{92.1})$
Gain $\texttt{+}10.3$ $(\texttt{+}7.2)$ $\texttt{+}11.5$ $(\texttt{+}8.4)$ $\texttt{+}11.2$ $(\texttt{+}8.0)$ $\texttt{+}10.4$ $(\texttt{+}8.0)$ $\texttt{+}12.3$ $(\texttt{+}9.4)$ $\texttt{+}15.2$ $(\texttt{+}11.0)$
Table 1: CIFAR-100 results. Results in terms of FAA []delimited-[][$\uparrow$][ ↑ ] and AIA []delimited-[][$\uparrow$][ ↑ ]. AIA is reported between parenthesis. Best results are highlighted in bold. Gain refers to the previous SOTA approach. The subscript FL indicates CL methods adapted for FCIL.
@@ 4.1 Settings
Following (Guo et al. 2024), we evaluated the proposed method on the CIFAR-100 (Krizhevsky, Hinton et al. 2009) and Tiny-ImageNet (Le and Yang 2015) datasets. Each one is partitioned into $10$ incremental tasks, resulting in $10$ and $20$ classes per task respectively. We distribute the data across $10$ clients, employing the widely adopted distribution-based (Li et al. 2022; Yurochkin et al. 2019) and quantity-based (Li et al. 2022; McMahan et al. 2017) label imbalance settings. The former partitions the data based on a Dirichlet distribution governed by a $\beta$ parameter, while the latter ensures that each client encounters exactly $\alpha$ classes within each task. We assess all methods across $6$ scenarios for each dataset. For quantity-based label imbalance settings, we set $\alpha\in\{2,4,6\}$ for CIFAR-100, and $\alpha\in\{4,8,12\}$ for Tiny-ImageNet. On both datasets, we conduct experiments with $\beta\in\{0.05,0.1,0.5\}$. For all experiments, we evaluate the centralized model using the global test set.
@@
We utilize a pre-trained ViT-B/16 as the backbone for HGP and all the compared methods. Specifically, we initialize the models with supervised pre-trained weights on ImageNet-21K (Ridnik et al. 2021) for CIFAR-100 and self-supervised pre-trained weights of DINO (Caron et al. 2021) for Tiny-ImageNet. This choice is grounded on mitigating potential data leakage when adapting the model for the latter dataset, which is a subset of ImageNet (Deng et al. 2009). For a complete overview of all the hyperparameters, refer to the supplementary material. Results are averaged across $3$ runs.

### v4

@@ Abstract
Federated Learning (FL) aims at unburdening the training of deep models by distributing computation across multiple devices (clients) while safeguarding data privacy. On top of that, Federated Continual Learning (FCL) also accounts for data distribution evolving over time, mirroring the dynamic nature of real-world environments. While previous studies have identified Catastrophic Forgetting and Client Drift as primary causes of performance degradation in FCL, we shed light on the importance of Incremental Bias and Federated Bias, which cause models to prioritize classes that are recently introduced or locally predominant, respectively. Our proposal constrains both biases in the last layer by efficiently fine-tuning a pre-trained backbone using learnable prompts, resulting in clients that produce less biased representations and more biased classifiers. Therefore, instead of solely relying on parameter aggregation, we leverage generative prototypes to effectively balance the predictions of the global model. Our proposed methodology significantly improves the current State Of The Art across six datasets, each including three different scenarios.
**footnotetext: Equal contribution
@@ 1 Introduction
The traditional paradigm in Deep Learning necessitates accessing large-scale datasets all at once, which hinders scalability and raises significant privacy concerns, especially when sensitive data is involved. Although distributing training across many devices could be an effective solution, there is still no effective mechanism for blending the resulting trained models into a single unified one. Federated Learning (FL) mcmahan2017communication addresses this challenge through a centralized server that coordinates distributed devices, aiming to create a single unified model while minimizing communication costs.
@@
Federated Class-Incremental Learning (FCIL) yoon2021federated ; dong2022federated ; zhang2023target takes a step further and couples distributed training with Online Learning, tolerating distribution shifts in the data over time. This presents new challenges, as deep models learning online (without relying on old examples) experience severe performance degradation due to Catastrophic Forgetting mccloskey1989catastrophic . In FCIL, the training process unfolds in tasks, each of which shifts the data distribution by introducing new categories. Each task is divided into communication rounds, wherein the local models train on their private data distribution. After local training, each client may transmit information to the orchestrator (server), which creates a global model and redistributes it to all clients. In the literature, some methodologies account for architectural heterogeneity (i.e., heterogeneous FL diao2021heterofl ; kim2022depthfl ; ilhan2023scalefl ), while others aim to enhance the performance of local models without necessarily converging to a global one (i.e., personalized FL collins2021exploiting ; ma2022layer ; oh2022fedbabu ). Instead, we follow the original FCIL setting as presented in dong2022federated , with the goal of training a single global model in a distributed way.
@@
When learning on a sequence of tasks, the model struggles the most at differentiating classes from distinct tasks, whereas it works well at separating those within the same one. Albeit one would intuitively link such behavior to Catastrophic Forgetting, it primarily occurs because tasks are learned separately, and some classes are never seen simultaneously kim2022theoretical . This results in a phenomenon known in the Incremental Learning literature as bias, which favors recently introduced classes wu2019large . We refer to this as Incremental Bias (IB). IB emerges because new classification heads are optimized independently, without concurrent access to previous classes. As a result, gradient updates are disproportionately influenced by the new classes, causing imbalance in the classifier’s output.
@@
Figure 1: Federated bias. Histogram of the clients’ responses computed on the global test set ($\beta=0.05$). Class indexes have been rearranged for improved visualization (left). The entropy of the response histograms, averaged on all clients, compared with FL performance (right).
@@
The authors of luo2021no observe a similar tendency in the Federated Learning scenario: since clients train exclusively on their local datasets, they exhibit a bias towards their local label distribution. We refer to this effect as Federated Bias (FB). In contrast to the well-known Client Drift gao2022feddc ; karimireddy2020scaffold ; zhao2018federated , which causes misalignment between the clients’ learned parameters, Federated Bias affects the clients’ responses. Specifically, FB induces clients’ outputs to diverge in different directions, mirroring the patterns of their local label distributions. Also, the strength of FB increases with growing heterogeneity, suggesting a correlation with declining performance (see Section 2). To relieve such an effect, we constrain FB to the last layer by leveraging a frozen pre-trained backbone and efficiently fine-tuning it via prompt learning li2021prefix . Ideally, prompting keeps the clients’ representations close to the pre-training optimum (hence, close to each other), thus minimizing their Federated Bias. In Section 4.3, we experimentally verify that prompting leads to reduced bias in the feature space w.r.t. fine-tuning all parameters. This confines the impact of FB to the last layer, providing the centralized server with less biased representations. On top of that, prompt-based methodologies i) have demonstrated SOTA results in Class-Incremental Learning wang2022learning ; wang2022dualprompt ; smith2023coda and ii) adapt only a small portion of the clients’ parameters, improving communication efficiency in distributed scenarios zhao2023fedprompt ; liu2023fedet .
@@
The authors of wu2019large ; zhang2023slca ; luo2021no address either IB or FB by fine-tuning the classification layer (where such biases are the most evident) on IID data samples. To meet the privacy requirements of FL, which prohibit transferring real data, we follow recent studies zhang2023slca ; luo2021no and leverage latent generative replay. Specifically, at each communication round, we alleviate both biases – previously enforced to the last layer by the adopted prompt-based fine-tuning – by rebalancing the global classifier on a dataset of generated representations. In contrast to other approaches relying on prototypes (i.e., the average feature vectors) to regularize clients’ training procedures tan2022fedproto ; guo2024federated , we propose to compute their covariance matrix and parameterize a Multivariate Gaussian distribution for each class-client combination. This forms a grid of $num\_classes\times num\_clients$ generative prototypes, which are sampled hierarchically (first by class, then by client) to generate new data points.
@@ 2 On Federated Bias
This section investigates how Federated Bias in clients’ responses is associated with performance degradation in Federated Learning. All experiments are conducted on the CIFAR-100 krizhevsky2009cifar dataset, where the data is heterogeneously distributed across $10$ clients under the commonly adopted distribution-based label imbalance setting li2022federated ; yurochkin2019bayesian .
To effectively show the presence of FB in the local models, we evaluate them at their most biased state: specifically, at the end of the local training, prior to any synchronization with the centralized server. In Figure 1 (left), we show the histogram of the responses given by two randomly selected clients, trained with $\beta=0.05$, on the global test set. Additionally, we compare them against a model trained conventionally on the global data distribution (referred to as Joint). It can be observed that clients’ predictions are significantly skewed (i.e., biased), mirroring their local label distribution.
@@ 3.1 Problem definition
Federated Class-Incremental Learning zhang2023target ; guo2024federated tackles a classification problem across $C$ classes, which are introduced sequentially over $T$ incremental tasks. For each task, the data is distributed in a non-IID manner among $M$ clients. Let $D^{t}$ be the global partition for task $t$, which is split among the $M$ clients, with $D_{m}^{t}$ being the local partition of client $m$ at task $t$. The training procedure of each task is divided into communication rounds, each consisting of a certain number of epochs. At the end of the local optimization, the clients synchronize with the server by exchanging their learnable parameters. The server aggregates these parameters into a global model and redistributes it to all clients, thus concluding the communication round. Since data from previous tasks is unavailable, client $m$ can only train on its dataset $D_{m}^{t}$ during task $t$. Let $f_{\theta_{m}}$ be the local model of client $m$, parameterized by $\theta_{m}$. The local objective is to minimize the loss function $\mathcal{L}$ with respect to the local dataset $D_{m}^{t}$, namely:
@@ 3.2 Hierarchical Generative Prototypes
Differently from other prompt-based CL approaches, which rely on a pool of task-generic wang2022learning or task-specific smith2023coda ; wang2022dualprompt prompts, we propose learning a single prompt shared across all tasks. This eliminates the need for prompt selection – which requires an additional forward pass – significantly accelerating both training and inference. Specifically, we instantiate two learnable vectors, $P_{k}$ and $P_{v}$, for each of the first $5$ transformer blocks, obtaining the prompt $\mathcal{P}=\{P_{k}^{j},P_{v}^{j}\;|\;j\in 1,\ldots,5\}$. Following prefix-tuning li2021prefix , $P_{k}^{j}$ and $P_{v}^{j}$ are respectively prepended to the keys and values of the $j^{th}$ Multi-head Self Attention layer and optimized jointly to minimize the loss function.
@@
For a given class $c$, the server receives $M$ generative prototypes, one from each client, and seeks to identify the single distribution that best aligns with all of them. To this aim, we take the distribution that minimizes the Jensen-Shannon Divergence (JSD) lin1991divergence . Considering multiple distributions $\mathcal{Q}=\{Q_{1},\ldots,Q_{M}\}$, with importance weights $\Pi=\{\pi_{1},\ldots,\pi_{M}\}$, their JSD is defined as:
@@
Notably, the optimal solution to this problem is to set $\tilde{Q}_{c}$ equal to the distribution $G$. Since, in our case, all distributions $\{Q_{1},\ldots,Q_{M}\}$ are Gaussians, the closest distribution $G$ is precisely defined as a Gaussian Mixture Model (GMM). Following a common practice in Federated Learning tan2022fedproto ; mcmahan2017communication , we assign the importance weights $\{\pi_{1,c},\ldots,\pi_{M,c}\}$ to the generative prototypes based on the number of samples from class $c$ observed by client $m$.
@@
Finally, the server redistributes the updated classifier along with the aggregated prompt to the clients, enabling them to start the next training round. We refer the reader to the supplementary material for a detailed analysis of the computational complexity of our approach.
@@ 4.1 Settings
Following salami2024closed , we evaluated the proposed method on six diverse datasets, spanning both in-distribution and out-of-distribution domains: CIFAR-100 krizhevsky2009cifar , ImageNet-R hendrycks2021many , ImageNet-A hendrycks2021natural , EuroSAT helber2018introducing , Cars-196 Krause_2013_ICCV_Workshops , and CUB-200 WahCUB_200_2011 . All datasets are partitioned into $10$ incremental tasks, except for EuroSAT, which is split into $5$ tasks. We distribute the data across $10$ clients using the widely adopted distribution-based label imbalance setting li2022federated ; yurochkin2019bayesian . This approach partitions the data according to a Dirichlet distribution controlled by a $\beta$ parameter. Additional results involving up to $100$ clients are provided in the supplementary material.
For each dataset, we experiment with three different values of $\beta$, selected based on the number of examples. Specifically, we use $\beta\in\{0.5,0.1,0.05\}$ for ImageNet-R and CIFAR-100, and $\beta\in\{1.0,0.5,0.2\}$ for EuroSAT, ImageNet-A, Cars-196 and CUB-200. This choice is imposed by the infeasibility of finding a valid split when either the data or the number of classes is limited. For all experiments, we evaluate the centralized model at the end of the training on the global test set.
@@
We use a pre-trained ViT-B/16 as the backbone for all compared methods, initializing the models with supervised pre-trained weights on ImageNet-21K ridnik2021imagenet . For a comprehensive overview of the hyperparameters, please refer to the supplementary material. The results are averaged over three runs on different seeds. We report the standard deviations in the suppl. material.